# Superconductivity, Anomalous Hall Effect, and Stripe Order in Rhombohedral Hexalayer Graphene

### v1

@@
Erin Morissette1    Peiyu Qin1    HT Wu1    K. Watanabe2    T. Taniguchi3    J.I.A. Li1 jia_li@brown.edu 1Department of Physics, Brown University, Providence, RI 02912, USA 2Research Center for Functional Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 3International Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(April 7, 2025)
We report the discovery of a unique superconducting phase in rhombohedral hexalayer graphene characterized by its simultaneous emergence with both the anomalous Hall effect and stripe charge order. The onset of stripe charge order is revealed through angle-resolved transport measurements, which show thermally activated insulating behavior along one axis and highly conductive transport along the orthogonal direction. Superconductivity develops exclusively along the high-conductivity axis, giving rise to a one-dimensional–like superconducting channel. This superconducting state exhibits first-order transitions under an out-of-plane magnetic field, consistent with a chiral order parameter that breaks time-reversal symmetry. Most remarkably, thermally driven superconducting transitions display pronounced hysteresis—an uncommon phenomenon that reflects the complex interplay among stripe formation, broken time-reversal symmetry, and superconductivity. Together, these results uncover a previously unidentified quantum phase: a chiral superconductor embedded within an anomalous Hall crystal.
@@
Figure 3: Hysteretic transitions driven by magnetic field and temperature temperature cycling. (a–b) Transport responses as a function of out-of-plane magnetic field $B$: (a) $R_{\perp}$ measured with current flowing parallel to the stripe direction; (b) $R_{\text{cross}}$ measured transverse to the stripes, with current flowing along $\phi=270^{\circ}$. (c–d) Temperature sweep of (c) $R_{\parallel}$ and (d) $R_{\perp}$ measured during both warming and cooling cycles with current applied perpendicular to the stripe orientation at $\phi=135^{\circ}$. (e–f) $R_{\parallel}$ versus $T$ measured during both warming and cooling cycles along (e) $\phi=225^{\circ}$ and (f) $270^{\circ}$, displaying prominent hysteresis in the superconducting transition.
Figure 4: Anisotropic superconducting transport across SC i. (a) Color-scale maps of the longitudinal resistance $R_{\parallel}$ measured with current flowing along the stripes ($\phi=225^{\circ}$, left panel) and transverse to the stripes ($\phi=135^{\circ}$, right panel), plotted as a function of carrier density $n$ and displacement field $D$. Dashed lines outline the boundary of the SC i pocket. White solid lines represent constant-$D$ line cuts analyzed in panels (b–c), while the white dash-dotted line indicates a diagonal cut examined in panel (d). (b–d) Horizontal line cuts at a constant displacement field (b) $D=0.93$, (c) $0.965$, and $0.97$ V/m, showing $R_{\parallel}$ (top) and $R_{\perp}$ (bottom). For clarity, the blue trace in top panels is multiplied by a factor of 5. (d) Longitudinal resistance $R_{\parallel}$ measured along the diagonal dashed line shown in panel (a), traversing both superconducting pockets SC i and SC ii.
@@
Chiral superconductivity has been extensively explored in the unique superfluid A-phase of 3He  Vollhardt and Wölfle (1990), which emerges at ultra-low temperatures around $2$ mK in the strong coupling regime. This remarkable state can be simultaneously viewed as a $p$-wave superconductor and an orbital ferromagnet  Ambegaokar et al. (1974); Walmsley and Golov (2012). To reveal this duality, 3He is confined to a slab geometry, where spatial confinement aligns the orbital magnetic moments perpendicular to the plane, producing a net magnetization at the onset of the superfluid A-phase. By investigating the transport response of moving electron bubble, the magnetic moment associated with the superfluid phase was directly detected via the intrinsic magnus force acting on electron bubbles Ikegami et al. (2013). This distinctive transport signature offers direct demonstration of chiral symmetry breaking in the order parameter of the superfluid A-phase of 3He.
More recently, new graphene-based platforms have emerged as fertile ground for exploring superconductivity intertwined with time-reversal symmetry breaking (TRSB). In particular, tetra- and pentalayer rhombohedral graphene exhibit superconductivity in the high-displacement-field regime that coexists with an anomalous Hall effect Han et al. (2024a); Parra-Martinez et al. (2025); Chen and Schrade (2025); Chou et al. (2024); Yang and Zhang (2024); Christos et al. (2025); Yoon et al. (2025). In these systems, orbital ferromagnetism appears to intertwine with the Cooper pairing instability, giving rise to magnetic-field–driven switching between superconducting states, establishing experimental support for a chiral superconducting state.
@@
Unlike the fluid phase of 3He, where chiral order emerges in relative isolation, the low-temperature phase diagrams of two-dimensional (2D) electron systems—particularly those governed by strong Coulomb interactions—often host a rich interplay of coexisting and competing orders. Notably, recent experiments have demonstrated that the emergence of TRSB in rhombohedral multilayer graphene is intimately tied to spontaneous rotational symmetry breaking Morissette et al. (2025). This connection raises the intriguing possibility that the chiral superconducting state in these systems features broken rotational symmetry—a hypothesis we explore in this work.
@@
To investigate this, we leverage recent advances in angle-resolved transport techniques Wu et al. (2017, 2020); Vafek (2023); Chichinadze et al. (2024); Zhang et al. (2024a); Morissette et al. (2025); Zhang et al. (2025), which provide powerful tools for characterizing how intertwined orders emerge and evolve alongside superconductivity. Applying these methods to rhombohedral hexalayer graphene, we uncover the prominent presence of a stripe charge order that coincides with both superconductivity and the anomalous Hall effect.
Central to our observation is the profound impact of stripe charge order on superconducting transport. Below the stripe formation temperature, charge transport becomes thermally activated and insulating along one principal axis, while remaining highly conductive along the orthogonal direction. Remarkably, at lower temperatures, the superconducting phase inherits this extreme anisotropy: dissipationless current is confined to the high-conductivity axis, forming a one-dimensional–like superconducting channel. By demonstrating the intertwined emergence of superconductivity, the anomalous Hall effect, and stripe charge order, our findings establish a highly unconventional setting for exploring topological superconductivity in low-dimensional confinements  Read and Green (2000); Qi and Zhang (2011); Nadj-Perge et al. (2014); Hua et al. (2024).
The low-temperature phase diagram of rhombohedral hexalayer graphene hosts three distinct superconducting pockets (Fig. 1a), denoted SC i, SC ii, and SC iii in Fig. 1b. Their superconducting transition is demonstrated by vanishing longitudinal resistance with decreasing temperature, as shown by the $R$$T$ curve in Fig. 1c Although all three pockets are situated near orbital ferromagnetic regimes, SC i stands out for its prominent coexistence with the anomalous Hall effect. Consequently, our investigation is focused on the unconventional behavior of SC i, while also examining the angular symmetry in SC ii and SC iii.
@@
We begin by examining the transport properties in the metallic regime above the superconducting transition temperature. The sample geometry and angle-resolved measurement setup, previously discussed in earlier works  Zhang et al. (2024a); Chichinadze et al. (2024); Morissette et al. (2025); Zhang et al. (2025), are also detailed in the Methods section. Fig. 1d presents the angle dependence of longitudinal and transverse resistance, $R_{\parallel}$ and $R_{\perp}$, measured at $n$ and $D$ values within the SC i regime. At $T=30$ K, the transport response appears nearly isotropic, but anisotropy becomes pronounced as the temperature is lowered. According to Fig. 1d, the principal axes of transport anisotropy, defined by the directions of maximum and minimum resistance, are aligned along $\phi=45^{\circ}$ and $135^{\circ}$, respectively.
@@
Figure 2a shows $R_{\parallel}$ as a function of temperature measured along different $\phi$. The maximum and minimum resistance are observed along the principal axes of transport anisotropy, revealing a divergence at low temperature that signals the emergence of extreme transport anisotropy.
To quantify the strength of anisotropy, we define parameter $\xi=(R_{\max}-R_{\min})/(R_{\max}+R_{\min})$, equivalent to $\Delta R/R_{0}$ in Eq. M1-2  Morissette et al. (2025); Zhang et al. (2025); Vafek (2023). The temperature dependence of $\xi$ follows a Curie-Weiss Law (dashed line in Fig. 2b), suggesting a thermodynamic transition driven by nematic instabilities  Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022). Notably, transport along the principal axis of maximum resistivity (minimum conductivity) exhibits activated insulating behavior (inset of Fig. 2b). A linear fit to the Arrhenius plot reveals an energy gap associated with this insulating behavior, $\Delta\approx 4.5$ K. In stark contrast, charge transport along the orthogonal axis remains highly conductive. Such extreme anisotropy, combined with the activated insulating behavior along one principal axis, is consistent with a stripe charge order, with the easy axis near $\phi=45^{\circ}$ (marked by a blue arrow in the inset of Fig. 2b).
@@
Most remarkably, as superconductivity emerges at low temperature, its transport response inherits the spatial anisotropy of the stripe charge order. This extreme anisotropy is manifested in the strong dependence in transition temperature $T_{c}$ on the relative orientation between current flow and the stripe orientation. The highest $T_{c}$ is observed to be around $1$ K when current is applied along the stripe orientation. In contrast, when the current flow is misaligned from the stripe orientation by $45^{\circ}$, as shown in the inset of Fig. 2a, $T_{c}$ is suppressed by almost a factor of 2. The extreme scenario arises when current is applied transverse to the stripe orientation along $\phi=135^{\circ}$. Here, the insulating behavior persists down to the base temperature of the dilution fridge ($T=30$ mK).
@@
Crucially, the extreme transport anisotropy we observe is robust against variations in measurement setup. Beyond the configurations shown in Fig. 2a, Fig. 2c displays transport responses with the current applied along $\phi=90^{\circ}$. In this arrangement, voltage contacts (A, B) that probe the direction perpendicular to the stripes show a low-temperature insulating response, while contacts (A, B) along the stripe direction exhibit diminishing resistance at low temperature. For simplicity, we will refer to this setup as the cross configuration. A similar observation arises under an alternative configuration uses different source–drain and voltage contacts, as illustrated in Fig. 2d: once again, transport measurement reveals an insulating behavior transverse to the stripes and dissipationless parallel to them.
@@
These observations confirm that the coexistence of insulating and superconducting responses originates from the stripe formation, rather than an artifact of specific contact geometry. Additional configurations in Fig. M9 and Fig. M10 further demonstrate that the same set of source-drain-voltage contacts can yield either an insulating or a superconducting response, depending on their orientation relative to the stripe order—ruling out poor electrical contacts as the cause.
A notable and perhaps defining characteristic of the SC i phase is the prevalence of hysteretic behaviors in transport responses. In the following, we systematically examine these hysteretic transitions, providing a unique window into the complex interplay between spontaneous valley polarization, stripe charge order, and superconductivity.
First, we demonstrate that the superconducting phase undergoes hysteretic transitions driven by an out-of-plane magnetic field ($B_{\perp}$), consistent with recent observations in rhombohedral tetra- and pentalayer graphene Han et al. (2024a). When current flows parallel to the stripe direction, Figures 3a–b reveal clear hysteretic switching between two superconducting ground states as $B_{\perp}$ is varied. This behavior is indicative of a first-order transition between degenerate order-parameter configurations, closely analogous to the A-phase of superfluid 3He—a prototypical chiral superconductor Ikegami et al. (2013). Remarkably, similar $B_{\perp}$-driven hysteretic switching is observed even when the current flows perpendicular to the stripe orientation. As shown in the cross configuration of Fig. 3b, hysteresis remains visible despite the transport response being highly resistive. Collectively, these findings establish that the anomalous Hall effect persists throughout both the superconducting and insulating regimes, unveiling a highly exotic scenario in which superconductivity is embedded within an anomalous Hall crystal.
@@
The presence of a crystalline phase necessitates a first-order melting transition, which manifests as hysteresis due to supercooling and superheating during thermal cycling. Consistent with this expectation, temperature-driven hysteresis is clearly observed when probing the insulating response with current applied perpendicular to the stripe direction. As shown in Figs. 3c–d, both $R_{\parallel}$ and $R_{\perp}$ measured along $\phi=135^{\circ}$ exhibit pronounced hysteresis loops during controlled warming and cooling cycles (see Methods for experimental details). Crucially, the shape and magnitude of these loops cannot be attributed to artifacts in temperature regulation. For example, $R_{\perp}$  exhibits a distinct resistance peak during cooling that is completely absent during warming, underscoring its intrinsic origin.
Even more strikingly, hysteretic behavior also emerges across the superconducting transition itself. As illustrated in Figs.3e–f, the superconducting transition temperature $T_{c}$, defined by the vanishing of $R_{\parallel}$, is significantly suppressed on cooling. Notably, the suppressed $T_{c}$ aligns closely with the resistance peak observed in the insulating response during the same cooling cycle (Fig.3d), revealing a compelling connection between the onset of superconductivity and the underlying stripe order.
@@
The observed hysteresis in the superconducting transition is highly unusual. First-order transitions in superconductors typically occur between distinct superconducting phases with different order parameters, as exemplified by the A- and B-phases of superfluid 3He Vollhardt and Wölfle (1990). In contrast, transitions between the superconducting and normal states are generally expected to be continuous and second-order. In our system, however, several plausible mechanisms could give rise to a first-order superconducting transition:
@@
(i) The melting of the stripe charge order is itself a first-order transition. Given the close proximity in temperature between stripe formation and the onset of superconductivity, their coupling could naturally result in a discontinuous superconducting transition with associated hysteresis.
@@
(ii) The superconducting transition temperature ($T_{c}$) is extremely sensitive to the relative orientation between the applied current and the stripe direction. In this scenario, the hysteresis arises from subtle changes in stripe orientation induced either by thermal cycling above $T_{\text{stripe}}$ or by current-direction switching during the measurement sequence.
@@
(iii) There may exist a temperature-driven transition between two distinct superconducting phases, analogous to the A- and B-phases of superfluid 3He. Supercooling and superheating across such a first-order transition would naturally lead to the observed pronounced hysteresis  Schiffer et al. (1992); Schiffer and Osheroff (1995); Tian et al. (2023).
@@
Taken together, our observations reveal a remarkable and intricate interplay among multiple emergent electronic orders. The magnetic-field–driven hysteretic switching between distinct superconducting ground states strongly supports the presence of a chiral superconducting phase Han et al. (2024a). Simultaneously, the coexistence of stripe charge order and the anomalous Hall effect offers compelling experimental evidence for an anomalous Hall crystal state Tan and Devakul (2024); Dong et al. (2024); Soejima et al. (2024). Moreover, the coexistence of superconductivity with stripe charge order naturally leads to a spatial modulation of the superconducting order parameter, positioning this system as a promising platform for realizing pair density wave states Agterberg et al. (2020).
Previous studies of rhombohedral pentalayer graphene aligned with hexagonal boron nitride (hBN) have also hinted at an anomalous Hall crystal phase Lu et al. (2024, 2025). Our observations of stripe order coexisting with the anomalous Hall effect within a similar region of the $n$$D$ phase diagram suggests that stripe formation may represent a universal ordering tendency in rhombohedral multilayer graphene under strong displacement fields.
Beyond the anomalous Hall crystal interpretation, the coexistence of stripe order and anomalous Hall transport might also be influenced by disorder-induced localization. While conventional transport measurements alone cannot definitively exclude disorder, our sequential measurements, with rotating current flow direction (see Methods), reveal that the stripe orientation can be reconfigured by applied current—a behavior more consistent with an interaction-driven crystalline order than with static disorder. The possibility of such an emergent crystal phase will undoubtedly motivate future experimental investigations.
The stripe order plays a crucial role shaping the properties of superconducting transport. For instance, an outstanding signature of the superconducting phase in rhombohedral tetra- and pentalayer graphene is its remarkable stability against a large out-of-plane magnetic field ($B_{\perp}$)  Han et al. (2024a). In rhombohedral hexalayer graphene, however, this stability exhibits a strong dependence on the orientation of the applied current relative to the stripe direction. When current flows along the stripes, $R_{\parallel}$ remains vanishingly small even at $B_{\perp}=0.5$ T (dark blue trace in Fig.4a), and the corresponding $R$$T$ curves display a sharp superconducting transition, persisting despite the substantial out-of-plane magnetic field.
In contrast, when current is applied along $\phi=270^{\circ}$—misaligned by approximately $45^{\circ}$ from the stripe direction—the application of $B_{\perp}=-0.3$ T leads to a significant increase in $R_{\parallel}$ at low temperatures. Nevertheless, the onset of superconductivity, as indicated by a sharp deviation from the normal-state behavior (highlighted by vertical dashed lines in Fig. 4c), remains largely unaffected by the magnetic field.
The observation of robust superconducting transport at $B=0$ along $\phi=270^{\circ}$ raises an intriguing question about the nature of inter-stripe coupling. As shown in the distinct $R$$T$ behavior in Figs. 4b–c, superconductivity first emerges along the stripe direction near $1$ K and remains robust against $B_{\perp}=-0.3$ T. These intra-stripe channels govern the initial onset of superconductivity. In contrast, the vanishing of $R_{\parallel}$ along $\phi=270^{\circ}$ at low temperatures relies on inter-stripe coupling between adjacent stripes. The application of $B_{\perp}$ primarily disrupts this inter-stripe coupling, leading to a finite $R_{\parallel}$ at low temperatures, while leaving the superconducting onset largely unaffected.
With further increase in $B_{\perp}$, superconductivity is completely suppressed at $B_{\perp}=-1$ T. In this regime, $R$$T$ curves exhibit insulating behavior at low temperatures, irrespective of current flow direction. Notably, the superconducting transition temperature ($T_{c}$) in rhombohedral hexalayer graphene is significantly higher than values reported for tetra- and pentalayer counterparts Han et al. (2024a). This enhanced $T_{c}$ may stem from stronger Coulomb interactions in the flatter bands of hexalayer graphene, as suggested by both theoretical predictions and experimental observations Zhang et al. (2024b); Muten et al. (2021); Pamuk et al. (2017).
The emergence of stripe order is a universal feature across the entire parameter space of the SC i phase. When current flows perpendicular to the stripes, a highly resistive transport response is observed (Fig. 4f), signaling the apparent absence of superconductivity in this configuration. This behavior stands in stark contrast to the vanishing $R_{\parallel}$ observed when current is applied parallel to the stripe orientation (Fig. 4e). The anisotropic nature of the SCi phase is further highlighted by line cuts taken at fixed displacement field $D$ (Fig. 4g; see also Fig. M7).
Even when current is applied transverse to the stripe direction, $R_{\perp}$ drops to zero within the SC i region (Fig. 3h). This observation is consistent with superconductivity developing exclusively along the stripes, which equalizes the potential along their direction, thus resulting in a vanishing $R_{\perp}$ despite a finite, resistive $R_{\parallel}$. Moreover, the extreme anisotropy appears to be unique to the SC i phase. Fig. M8 shows the transport response measured along a diagonal line crossing both the SC i and SC ii phases. While $R_{\parallel}$ exhibits a pronounced discrepancy between the directions parallel and perpendicular to the stripe order in SC i, this discrepancy vanishes in SC ii. The transport anisotropy observed in SC ii is discussed in more detail in the Methods section.
Overall, we employed angle-resolved transport measurements to reveal an unusual superconducting phase embedded in an anomalous Hall crystal. Although this unconventional coexistence has not been explicitly predicted, several theoretical studies have examined Coulomb-driven instabilities involving valley degrees of freedom, where the simultaneous breaking of time-reversal and rotational symmetries promotes spontaneous condensation in momentum space Dong et al. (2023); Jung et al. (2015); Huang et al. (2023); Mandal and Fernandes (2023); Parra-Martinez et al. (2025). Such momentum-space condensation has been widely studied in various multilayer graphene systems Zhang et al. (2024a); Lin et al. (2023), and was recently identified in the multiferroic phase of rhombohedral hexalayer graphene near zero displacement field ($D\approx 0$Morissette et al. (2025).

### v2

@@
Erin Morissette1    Peiyu Qin1    Hai-Tian Wu1    Naiyuan J. Zhang1    K. Watanabe2    T. Taniguchi3    J.I.A. Li1 jia_li@brown.edu 1Department of Physics, Brown University, Providence, RI 02912, USA 2Research Center for Functional Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 3International Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(April 16, 2025)
Unconventional superconducting phases are distinguished by broken symmetries in their order parameters. Here, we report the discovery of an exotic superconducting phase in rhombohedral hexalayer graphene at high displacement fields, marked by its coexistence with both a stripe charge order and the anomalous Hall effect. In angle-resolved transport measurements, the onset of stripe order is manifested as thermally activated insulating behavior along one axis, with highly conductive transport along the orthogonal direction. Upon cooling, superconductivity emerges exclusively along the high-conductivity axis, forming one-dimensional–like superconducting channels. This anisotropic superconducting phase exhibits multiple first-order hysteretic transitions. Pronounced thermal hysteresis-between warming and cooling across the superconducting transition-suggests a melting transition of the underlying stripe phase. Additionally, magnetic-field–induced switching—similar to phenomena reported in tetra- and pentalayer graphene—supports the chiral nature of the superconducting state. Together, these findings identify a previously unrecognized quantum phase: a chiral superconductor embedded within a stripy Hall crystal.
@@
Figure 3: Temperature-driven hysteretic transitions and stripe melting. (c–d) Temperature sweep of (c) $R_{\parallel}$ and (d) $R_{\perp}$ measured during both warming and cooling cycles with current applied perpendicular to the stripe orientation at $\phi=135^{\circ}$. (e–f) $R_{\parallel}$ versus $T$ measured during both warming and cooling cycles along (e) $\phi=225^{\circ}$ and (f) $270^{\circ}$, displaying prominent hysteresis in the superconducting transition. Measurements are performed at $D=958$ mV/nm and $n=8\times 10^{12}$ cm-2.
Figure 4: Magnetic-field-driven hysteretic transitions and chiral superconductivity. (a–f) $R_{\parallel}$ (top panels) and $R_{\perp}$ (bottom panels) as a function of out-of-plane magnetic field $B$ measured with current flowing parallel to the stripe direction at (a) $T=0.11$ K, (b) $0.48$ K, (c) $0.81$ K, (d) $1.40$ K, and (e) $1.80$ K. Measurements in (a–f) are performed at $D=930$ mV/nm and $n=8\times 10^{12}$ cm-2. (g-i) $R_{\parallel}$ as a function of $T$ measured at different values of out-of-plane $B$-field and current flow directions.
@@
The superfluid A-phase of 3He, which emerges at ultra-low temperatures around 2 mK in the strong-coupling regime, provides a prototypical example of chiral superconductivity Vollhardt and Wölfle (1990). In this remarkable state, Cooper pairing occurs with non-zero angular momentum, allowing the phase to be simultaneously understood as a $p$-wave superconductor and an orbital ferromagnet Ambegaokar et al. (1974); Walmsley and Golov (2012). This duality is manifested in a net magnetization associated with the angular momentum of the Cooper pairs. When confined to a slab geometry, spatial confinement aligns the orbital magnetic moments perpendicular to the plane, producing two degenerate configurations with magnetic moments pointing either up or down Ikegami et al. (2013). Most remarkably, applying an out-of-plane magnetic field drives a transition between these configurations by reversing the direction of the magnetic moment—a process that manifests in the transport response of moving electron bubbles. This distinctive transport signature provides direct evidence of the chiral nature of the superfluid A-phase of 3He.
Despite decades of effort to identify a chiral superconductor in a solid-state platform, the superfluid A-phase of 3He remained the only known example until recently. A promising candidate has now emerged in the superconducting phases of tetra- and pentalayer rhombohedral graphene Han et al. (2024a); Parra-Martinez et al. (2025); Chen and Schrade (2025); Chou et al. (2024); Yang and Zhang (2024); Christos et al. (2025); Yoon et al. (2025). In these systems, the Cooper pairing instability appears to be intimately linked to orbital ferromagnetism. The resulting superconducting phase exhibits magnetic-field–driven switching between distinct ground states—closely resembling the hallmark behavior of the 3He A-phase Han et al. (2024a).
@@
Unlike the fluid phase of 3He, where chiral order emerges in relative isolation, the low-temperature phase diagrams of two-dimensional (2D) electron systems—particularly those governed by strong Coulomb interactions—often host a rich interplay of coexisting and competing orders Fradkin et al. (2015); Keimer and Moore (2017); Fernandes et al. (2019); Nuckolls and Yazdani (2024); Wang et al. (2023). Among these, multilayer graphene with rhombohedral stacking has emerged as a particularly fertile platform, where the emergence of superconductivity is intertwined with a broad spectrum of electronic orders Zhou et al. (2021a, b, 2022); Han et al. (2024a); Choi et al. (2025). Notably, recent experiments have revealed that the onset of time-reversal symmetry breaking (TRSB) in rhombohedral multilayer graphene is intimately linked to spontaneous rotational symmetry breaking Morissette et al. (2025). This observation raises the compelling possibility that the chiral superconducting state in these systems intrinsically features broken rotational symmetry—a hypothesis we investigate in this work.
@@
To investigate this possibility, we leverage recent advances in angle-resolved transport techniques  Wu et al. (2017, 2020); Vafek (2023); Chichinadze et al. (2024); Zhang et al. (2024a); Morissette et al. (2025); Zhang et al. (2025), which offer powerful tools for probing how intertwined electronic orders emerge and evolve alongside superconductivity. Applying these methods to rhombohedral hexalayer graphene, we uncover a prominent stripe charge order that coexists with both superconductivity and the anomalous Hall effect.
The low-temperature phase diagram of rhombohedral hexalayer graphene hosts three distinct superconducting pockets, as shown in Fig. 1a, labeled SCi, SCii, and SCiii in Fig. 1b. Their superconducting transitions are evident in the $R$$T$ characteristics displayed in Fig. 1c. While all three pockets are proximate to regimes of orbital ferromagnetism, SCi emerges from a valley-polarized metallic phase and uniquely exhibits a pronounced coexistence with stripe charge order (see Fig. M1 and Fig. M2). Accordingly, our study focuses on the unconventional properties of the SCi phase. For completeness, we also comment on the angular transport behavior observed in SCii and SCiii in the Methods section.
@@
We begin by examining the transport properties in the metallic regime above the superconducting transition temperature. The sample geometry and angle-resolved measurement setup have been described in detail in previous studies Zhang et al. (2024a); Chichinadze et al. (2024); Morissette et al. (2025); Zhang et al. (2025). Figure 1d presents the angular dependence of both the longitudinal and transverse resistances, $R_{\parallel}$ and $R_{\perp}$, measured at carrier density $n$ and displacement field $D$ values corresponding to the SCi regime. At $T=30$,K, the transport response is nearly isotropic, but as the temperature decreases, a pronounced anisotropy emerges. As shown in Fig.1d, the principal axes of transport anisotropy—defined by the directions of maximum and minimum resistance—are aligned along $\phi=45^{\circ}$ and $135^{\circ}$, respectively.
@@
Figure 2a shows $R_{\parallel}$ as a function of temperature, measured at various angles $\phi$. The maximum and minimum resistance values, $R_{\max}$ and $R_{\min}$, are observed when current is applied along the principal axes of the transport anisotropy. The increasing divergence between $R_{\max}$ and $R_{\min}$ at low temperatures signals the emergence of extreme transport anisotropy.
We define the anisotropy parameter as $\xi=(R_{\max}-R_{\min})/(R_{\max}+R_{\min})$, which is equivalent to $\Delta R/R_{0}$ in Eq. M1–M2 Morissette et al. (2025); Zhang et al. (2025); Vafek (2023). The dashed line in Fig. 2b represents the best fit of the temperature dependence of $\xi$ to the Curie–Weiss law. The excellent agreement between the experimental data and the Curie–Weiss fit indicates a thermodynamic transition driven by nematic instabilities Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022).
Notably, transport along the principal axis of maximum resistivity (i.e., minimum conductivity) exhibits thermally activated insulating behavior, as shown in the inset of Fig. 2b. A linear fit to the Arrhenius plot yields an energy gap of approximately $\Delta\approx 4.5$ K. In stark contrast, charge transport along the orthogonal axis remains highly conductive. The emergence of activated behavior along only one direction is an unprecedented observation. It points to the formation of an exotic electronic order that goes beyond conventional nematicity. We propose that our findings arise from the formation of a stripe charge order at low temperature  Fogler et al. (1996); Koulakov et al. (1996); Lilly et al. (1999a, b); Du et al. (1999); Pan et al. (1999), with the easy axis aligned near $\phi=45^{\circ}$ (indicated by the blue arrow in the inset of Fig. 2b).
@@
Central to our observation is the profound impact of stripe charge order on superconducting transport. As superconductivity emerges at low temperatures, its transport response inherits the spatial anisotropy imposed by the stripe order. When current is applied along the stripe orientation, the superconducting transition temperature $T_{c}$ reaches a maximum near 1 K. In contrast, when the current is misaligned by $45^{\circ}$—as shown in the inset of Fig. 2a—$T_{c}$ is suppressed by nearly a factor of two. A more dramatic suppression occurs when current is applied transverse to the stripe orientation, along $\phi=135^{\circ}$, where insulating behavior persists down to the base temperature of the dilution refrigerator ($T=30$ mK). This extreme anisotropy is reminiscent of quantum Hall stripe phases Lilly et al. (1999a, b); Du et al. (1999); Pan et al. (1999), though distinguished here by the emergence of superconductivity along the stripe direction.
@@
Crucially, the extreme transport anisotropy we observe is robust against variations in measurement configuration. For example, Fig. 2c shows an alternative setup with current applied along $\phi=90^{\circ}$. In this arrangement, voltage probes positioned perpendicular to the stripe direction (A, B) display a low-temperature insulating response, while probes aligned along the stripes (A, B) exhibit diminishing resistance as temperature decreases. For simplicity, we refer to this setup as the cross configuration. A similar observation is made using different source–drain and voltage contact pairs, as shown in Fig. 2d: once again, the transport response reveals insulating behavior transverse to the stripes and dissipationless behavior along them.
@@
Additional configurations in Fig. M10 and Fig. M11 further demonstrate that whether the response is insulating or superconducting is solely determined by the orientation relative to the stripe order—ruling out poor electrical contacts as the cause. These observations confirm that the coexistence of insulating and superconducting responses originates from the stripe formation, rather than from artifacts related to specific contact geometry.
In the following, we report a rich array of hysteretic behaviors in transport measurements that underscore the remarkable coexistence of superconductivity, stripe charge order, and the anomalous Hall effect.
@@
Most notably, the emergence of stripe order is accompanied by a crystalline phase transition, evidenced by a first-order melting process with pronounced supercooling and superheating. This transition manifests as temperature-driven hysteresis in the insulating response when current is applied perpendicular to the stripe direction. As shown in Figs. 3c–d, both $R_{\parallel}$ and $R_{\perp}$ measured along $\phi=135^{\circ}$ exhibit clear hysteresis loops during controlled warming and cooling (see Methods for details). Importantly, the magnitude and asymmetry of these loops—such as the resistance peak in $R_{\perp}$ observed only during cooling—offer crucial confirmation for their intrinsic origin and rule out experimental artifacts.
The hysteretic $R$$T$ curves also offer insight into the characteristic temperature associated with stripe formation. As shown in Figs. 3a–b, we define the stripe melting temperature ($T_{\text{stripe}}$) as the onset of temperature-driven hysteresis in the insulating response measured perpendicular to the stripe direction. In this context, the resistance cusp observed in $R_{\parallel}$ (Fig. 3a) reflects a thermodynamic transition driven by nematic instability, without involving crystalline order.
Remarkably, $T_{\text{stripe}}$ coincides with the onset of superconductivity when current flows along the stripe direction (Fig. 3c), highlighting a direct link between superconductivity and stripe formation.
Even more striking is the observation of hysteresis across the superconducting transition itself. As shown in Figs. 3c–d, the superconducting transition—defined by the vanishing of $R_{\parallel}$—occurs at a significantly lower temperature during cooling, indicative of a first-order transition.
@@
The presence of hysteresis in the superconducting transition is highly unusual. First-order transitions in superconductors typically occur between distinct superconducting phases with different order parameters, such as the A- and B-phases of superfluid ${}^{3}\mathrm{He}$ Vollhardt and Wölfle (1990). In contrast, transitions between superconducting and normal states are generally expected to be continuous and second-order. In our system, however, several plausible mechanisms may lead to a first-order superconducting transition:
  1. (i)
@@
    The melting of the stripe charge order is itself a first-order transition. Therefore, the hysteresis observed in the superconducting transition may naturally result from the supercooling and superheating associated with stripe melting.
  2. (ii)
@@
    The superconducting transition temperature ($T_{c}$) is extremely sensitive to the relative orientation between the applied current and the stripe direction. In this scenario, the hysteresis arises from subtle changes in stripe orientation induced either by thermal cycling above $T_{\text{stripe}}$ or by current-direction switching during the measurement sequence.
  3. (iii)
@@
    There may exist a temperature-driven transition between two distinct superconducting phases, analogous to the A- and B-phases of superfluid 3He. Supercooling and superheating across such a first-order transition would naturally lead to the observed pronounced hysteresis  Schiffer et al. (1992); Schiffer and Osheroff (1995); Tian et al. (2023).
@@
Next, we show that the superconducting phase undergoes hysteretic transitions driven by an out-of-plane magnetic field ($B_{\perp}$), indicative of a first-order transition between degenerate order-parameter configurations—closely analogous to the A-phase of superfluid 3He, a prototypical chiral superconductor Ikegami et al. (2013). Figures 4a–c present superconducting transport responses with current flowing parallel to the stripe direction, measured as $B_{\perp}$ is swept back and forth at different temperatures. Both $R_{\parallel}$ and $R_{\perp}$ remain vanishingly small across most of the $B$ range, while field-driven reversals of the underlying valley polarization are marked by prominent peaks in $R_{\perp}$.
Such magnetic-field–driven switching is consistent with recent observations in rhombohedral tetra- and pentalayer graphene Han et al. (2024a), and offers strong support for the chiral nature of the superconducting phase.
While the field-driven reversal of valley polarization manifests as pronounced peaks in $R_{\perp}$, $R_{\parallel}$ remains largely featureless below the stripe formation temperature ($T<0.8$ K). It has been proposed that a resistive signal in the superconducting phase arises when a domain boundary lies between voltage probes Han et al. (2024a). In this context, the contrasting behavior between $R_{\parallel}$ and $R_{\perp}$ may reflect a tendency for domain boundaries to align preferentially along the stripe direction. As temperature increases, field-driven reversal becomes visible in both $R_{\parallel}$ and $R_{\perp}$ around $T=0.8$ K, consistent with the stripe melting temperature identified in Fig. 3. In this regime, domain boundaries are no longer pinned to the stripe orientation.
The magnetic-field–driven hysteresis loop persists up to $T=1.4$ K, indicating that spontaneous valley polarization is a more robust order than both superconductivity and stripe formation. Intriguingly, the onset of the anomalous Hall effect coincides with the nematic transition temperature $T_{\text{nem}}$ (Fig. 3). This connection is further evidenced by the overlapping phase space where extreme transport anisotropy and the anomalous Hall effect coexist, as shown in Fig. M1. A similar correlation has been observed in the multiferroic phase of rhombohedral hexalayer graphene near zero displacement field ($D\approx 0$Morissette et al. (2025). These observations point to a potential common origin for time-reversal and rotational symmetry breaking.